OpenStudy (anonymous):

solve the exponential equation express the solution set in terms of natural logarithms 2^x+8=3

OpenStudy (anonymous):

\[2^{x+8}=3\]

OpenStudy (isaiah.feynman):

Well, you should apply the natural logarithm to both sides.

OpenStudy (dumbcow):

hint: \[\ln a^x = x \ln a\]

OpenStudy (anonymous):

still need more steps to figure it out

OpenStudy (isaiah.feynman):

\[\ln 2^{x+8} = \ln 3\]

OpenStudy (isaiah.feynman):

Can you continue?

OpenStudy (anonymous):

no sorry still lost

OpenStudy (isaiah.feynman):

Okay, you know that the logarithm of a power makes the power cones down? so the x+8 will come down in front of the ln2\[(x+8) \ln2 = \ln 3\]

OpenStudy (anonymous):

yes ok I remember that

OpenStudy (isaiah.feynman):

Can you solve for x?

OpenStudy (anonymous):

do I distribute ln 2 to x+8

OpenStudy (isaiah.feynman):

Nope, that would be a little tedious. Divide both sides by ln 2.

OpenStudy (unklerhaukus):

...natural log? why not log to base 2 \[2^{x+8}=3\\ \log_2(2^{x+8})=\log_2(3)\\ (x+8)\log_2(2)=\log_2(3)\\ x+8=\log_2(3)\\ x = ... \]

zepdrix (zepdrix):

@UnkleRhaukus Because the question requires the answer to be in terms of the `natural log`. So using log base 2 would require a change of base later on, seems tedious.

OpenStudy (unklerhaukus):

hmm, you are right, it is important to read the question properly,

OpenStudy (isaiah.feynman):

@UnkleRhaukus I would have used base 2.

OpenStudy (anonymous):

have to go I will workm on this later today